chocolateluva77:

the perimeter of a rectangle is 66m. one side is 15 m longer than the double other side. setup and solve a system of equations to find the dimensions of the rectangle. pls explain how u got the steps. ty.

chocolateluva77:

*66 m

surjithayer:

let one side=x other side=2x+15 Perimeter=2[(2x+15)+x]=2(2x+15+x)=2(3x+15)=6x+30 6x+30=66 6x=66-30=36 x=36/6=? this is one side . find other side by substituting the value of x in 2x+15

chocolateluva77:

what is the value of x?

surjithayer:

36/6=6

chocolateluva77:

do i substitute the value of x into the equation 2[(2x+15)+x]=2(2x+15+x)=2(3x+15)=6x+30 to get the answer?

surjithayer:

one side=x=36/6=6 m other side=2x+15=2(6)+15=12+15=27 m

chocolateluva77:

okay ty

surjithayer:

yw

brianagatica14:

@surjithayer wrote:
let one side=x other side=2x+15 Perimeter=2[(2x+15)+x]=2(2x+15+x)=2(3x+15)=6x+30 6x+30=66 6x=66-30=36 x=36/6=? this is one side . find other side by substituting the value of x in 2x+15
you deserve best response